OpenStudy (anonymous):

How is tan (pi/4)=1?

OpenStudy (anonymous):

well you know your unit circle, right? here's an example: http://mathprecalculus.wikispaces.com/Chapter+5.1+Unit+Circle+Approach pi/4 is 45 degrees the coordinating pair is \[\sqrt{2}/2 , \sqrt{2}/2\] (cosine, sine) and tangent is sine/cosine SO tan of pi/4 = \[(\sqrt{2}/2)/(\sqrt{2}/2)\] which is 1

OpenStudy (anonymous):

@WinterRaine I get it, thanks.

OpenStudy (anonymous):

your welcome!